Effect of Transactions on Current Position Analysis

Data pertaining to the current position of Newlan Company are as follows:

Cash $284,900
Temporary investments 146,700
Accounts and notes receivable (net) 431,600
Inventories 315,400
Prepaid expenses 16,600
Accounts payable 116,200
Notes payable (short-term) 166,000
Accrued expenses 49,800

Instructions:

1. Compute (a) the working capital, (b) the current ratio, and (c) the quick ratio. Round the current ratio and the quick ratio to one decimal place.

Working capital $
Current ratio
Quick ratio

2. Compute the working capital, the current ratio, and the quick ratio after each of the following transactions, and record the results in the appropriate columns. Consider each transaction separately and assume that only that transaction affects the data given above. Format working capital as whole dollars. Round the current ratio and the quick ratio to one decimal place.

Transaction Working Capital Current Ratio Quick Ratio
a. Sold temporary investments for cash at no gain or loss, $37,000. $
b. Paid accounts payable, $66,000. $
c. Purchased goods on account, $42,000. $
d. Paid notes payable, $83,000. $
e. Declared a cash dividend, $66,000. $
f. Declared a stock dividend on common stock, $20,000. $
g. Borrowed cash from bank on a long-term note, $166,000. $
h. Received cash on account, $56,500. $
i. Issued additional shares of stock for cash, $332,000. $
j. Paid cash for prepaid expenses, $33,200. $
Answer #1

Current assets = cash + temporary investments + accounts and notes receivables + inventory + Prepaid expenses

Current assets = 284,900 + 146700 + 431600 + 315400 + 16600 = 1,195,200

Current liabilities = Accounts payable + Notes payable + Accrued expenses

Current liabilities = 116200 + 166000 + 49800 = 332,000

working capital = current assets - current liabilities = 1,195,200 - 332,000 = 863,200

answer 1

formula working answer
Working capital current assets - current liabilities 1195200 - 332000 $863200
Current ratio current assets / current liabilities 1195200 / 332000 3.6
Quick ratio current assets - inventory -prepaid exp / current liabilities [1195200-315400-16600] / 332000 2.6

answer 2

Transaction [ effects] Working Capital Current Ratio Quick Ratio
a. cash increased and investment decreased same amount so no change $863200 3.6 2.6
b.cash decreased and decreased accounts payable, $66,000. $863200 1129200/266000 = 4.2 860200/266000= 3.2
c. inventory and accounts payable increased $42,000. 1237200-374000= 863200 1237200/374000 = 3.3 863200/374000 = 2.3
d. decreased cash and notes payable, $83,000. 1112200-249000=863200 1112200/249000 = 4.5 780200/249000=3.1
e. current liabilities increases by, $66,000. $1195200 - 398000= 797200 1195200/398000=3 863200/398000=2.2
f. no changes $863200 3.6 2.6
g. increases cash by, $166,000. 1361200-332000= 1029200 1361200/332000= 4.1 1029200/332000= 3.1
h. No changes $863200 3.6 2.6
i. Increased cash, $332,000. 1527200-332000= 1195200 1527200/332200=4.6 1195200/332000= 3.6
j. no changes $863000 3.6 2.5